Please help solve this problem and explain it to me .. Thank You.
Y
________ = 3
68

Answer by checkley71(8403)

Y/63=3/1 OR CROSS MULTIPLYING WE GET Y=3*63 OR Y=189
PROOF 189/63=3 OR 3=3